A 4 ft piece of PVC pipe is cut in three pieces. The longest piece is 12 in. longer than twice the smallest piece. The middle piece is 6 in. less than the longest piece. How long is each piece?
Answer by RAY100(1637)

Let a,,b,,c,,,be the pieces,,,a,, largest,,,c,,,smallest,,,,in inches
.
a+b+c=4ft = 48 inches
.
a=2c+12,,,,,b= a-6,,,,,,,substituting
.
(2c+12) +(a-6) +c =48,,,,,,but a=(2c+12)
.
(2c+12+ +{(2c+12)-6} +c = 48
.
2c+12 +2c+12 -6 +c = 48
.
5c = 30
.
c=6,,,,a=2(6)+12= 24,,,,b=(26-6)= 18
.
.
check,,,,,a+b+c=48,,,,,24+18+6=48,,,,ok
.